OpenStudy (anonymous):

what is √x+8=x-4 A.8 B.-1 C.-4,4 D. no real solution

OpenStudy (anonymous):

Is the sq root over the x or x+8

OpenStudy (anonymous):

x+8

OpenStudy (anonymous):

I get x = 16....not one of your choices. OK that makes sense

OpenStudy (anonymous):

I get A. Do you want to see how to do this?

OpenStudy (anonymous):

yes

OpenStudy (luigi0210):

You really need to stop giving away answers ._.

OpenStudy (raden):

√(x+8)=x-4 square both sides, get x+8 = x^2 - 8x + 16 simplify by combine similar terms, get x^2 - 9x + 8 = 0 factor out above, (x-1)(x-8) = 0 setting each factor equals zero x-1=0 then x = 1 x-8=0 then x=8 we must recheck again, if x=1 and x=8 is right √(x+8)=x-4 put x=1, get √(1+8) = 1-4 3 = -3 this is false √(x+8)=x-4 put x=8 √(8+8)=8-4 4 = 4 nah, this is true so, only x=8 as the solution

OpenStudy (raden):

does that make sense, @karlisa1604 ?

OpenStudy (anonymous):

Yes. Thank you

OpenStudy (raden):

you're welcome